Need some help please. Huge final due soon.
Solve the following systems of equations
x+9y=8 (1)
x=5-9y (2)

Answer by algebrahouse.com(1659)

Need some help please. Huge final due soon.
Solve the following systems of equations

x + 9y = 8
x = 5 - 9y

Substitute (5 - 9y), in for x, into top equation

5 - 9y + 9y = 8 {substituted 5 - 9y, in for x, into top equation}
5 = 8 {combined like terms}

The variable disappeared.....most people now panic

When the variable disappears and you get a statement that is never true,
for example 5 = 8, then there is no solution.

*Also note*
When the variable disappears and you get a statement that is always true,
for example 3 = 3, then there are infinite solutions

The answer to this one is "no solution"
www.algebrahouse.com